# Event Spectroscopy: Event-based Multispectral and Depth Sensing using Structured Light

### v1

@@ Abstract
Uncrewed aerial vehicles (UAVs) are increasingly deployed in forest environments for tasks such as environmental monitoring and search and rescue, which require safe navigation through dense foliage and precise data collection. Traditional sensing approaches, including passive multispectral and RGB imaging, suffer from latency, poor depth resolution, and strong dependence on ambient light—especially under forest canopies. In this work, we present a novel event spectroscopy system that simultaneously enables high-resolution, low-latency depth reconstruction and multispectral imaging using a single sensor. Depth is reconstructed using structured light, and by modulating the wavelength of the projected structured light, our system captures spectral information in controlled bands between $650$ nm and $850$ nm. We demonstrate up to 60% improvement in RMSE over commercial depth sensors and validate the spectral accuracy against a reference spectrometer and commercial multispectral cameras, demonstrating comparable performance. A portable version limited to RGB (3 wavelengths) is used to collect real-world depth and spectral data from a Masoala Rainforest. We demonstrate the use of this prototype for color image reconstruction and material differentiation between leaves and branches using spectral and depth data. Our results show that adding depth (available at no extra effort with our setup) to material differentiation improves the accuracy by over 30% compared to color-only method. Our system, tested in both lab and real-world rainforest environments, shows strong performance in depth estimation, RGB reconstruction, and material differentiation—paving the way for lightweight, integrated, and robust UAV perception and data collection in complex natural environments.
@@ I Introduction
Figure 1: Event Spectroscopy: We propose an all-in-one, ambient-light independent solution for depth sensing, color image reconstruction and multispectral sensing. Usable for instance, for an uncrewed aerial vehicle (UAV) navigating in forest environments (bottom left). A) Our setup consisting of an event camera and a projector as an illumination source B) Generated depth of scene acquired using structured light C) Spectral image reconstructed using events D) Material segmentation of leaves using spectral and depth data.
There is an increased demand for uncrewed aerial vehicles (UAVs) flying in forests [1, 2, 3], both for environmental monitoring [4, 5, 6], and for search and rescue applications [7]. Vision is essential for obstacle sensing and data collection to enable these applications. For instance, environmental monitoring tasks such as sensor deployment [5, 8] or sample collection [4, 6, 9] require UAVs to fly in close proximity to tree branches and foliage, where fast-moving, thin structures must be detected under dynamic and often challenging lighting conditions. Flying in close proximity to trees makes interaction with foliage probable, and while it has been shown that UAVs can push aside flexible twigs and leaves while avoiding thicker branches[10], proper sensing to differentiate between branches and foliage is needed[11]. This requires not only low latency and high resolution depth sensing of thin and fine structures, but also accurate differentiation between woody branches and foliage.
@@
In this work, we propose a single event-based structured light solution which can simultaneously deliver high-resolution and low latency depth reconstruction with integrated multispectral sensing. The integrated system utilizes an event camera with structured light to reconstruct high-fidelity depth with low latency. By changing the wavelength of the light used for structured light we can not only reconstruct RGB images, but also perform multispectral sensing by projecting the desired wavelength - with lower latency than available multispectral sensors. First, we validate the quality of our depth reconstruction by demonstrating up to 60% average improvement in the RMSE over other commercial off-the-shelf depth sensors. Next, we demonstrate event-based multispectral sensing using wavelengths between 650 nm to 850 nm in a lab setting and compare our results to a commercial multispectral sensor and a ground truth spectrometer. Furthermore, we demonstrate RGB color image reconstruction using a portable version limited to RGB light projection, as well as material differentiation utilizing both the depth and spectral data on real-world data from a Masoala Rainforest.
@@ II Related Works
Our solution provides event-based multispectral and depth sensing. This requires both spectral imaging as well as event-based luminance recovery and geometry estimation. In the following section, current multispectral sensing methodologies are showcased, including how our system relates to existing systems. To provide depth sensing and to relate event measurements to scene luminance, existing approaches to event-based luminance recovery and geometry estimation are examined, and our method presented.
@@ II-A Spectral Imaging
Spectral imaging captures images in multiple wavelength bands, with established applications in precision agriculture, astronomy, and beyond. The resulting $(X,Y,\lambda)$ data cube contains both spatial $(X,Y)$ information and spectral information, $\lambda$, per wavelength. Data capture methods can be divided into two main groups: scanning and snapshot methods. Snapshot methods capture the full data cube during a single integration period of the sensor, whereas scanning methods achieve the same over multiple periods. Among scanning methods, tuned filters represent a popular option. These include a rotating filter wheel, an electro-mechanical Fabry-Ferot filter [14, 15], liquid-crystals [16], or acousto-optic tunable filters [17].
@@
Recent developments in compressive sensing like CASSI [20] use a two-dimensional patterned grating to reconstruct all images of all wavelengths in a single shot. While the computational complexity required to decode the image presents a major bottleneck, improvements have made this method viable for high-resolution imaging in both spatial and spectral dimensions [20, 21, 22].
In this work, we employ an inverse variant of a turned filter. Instead of filtering the light entering the imaging sensor, we illuminate the scene with light of a specific wavelength band, and measure the resulting change in reflected light off the scene with respect to ambient illumination.
II-B Event-based Luminance Recovery and Geometry Estimation
@@
Event-based luminance recovery is an evolving field of research that is undergoing significant exploration. As events naturally compress visual information, estimating the absolute intensity solely from events is a challenging task. Prior methods demonstrate reconstruction of a grayscale image up to an unknown intensity value through the use of data-driven priors [23]. Color intensity can also be recovered by using color event cameras [24], or through setups with multiple cameras and appropriate color filters [25]. Multispectral sensing for face recognition through the addition of a single infrared filter was shown in [26]. In [27], the illumination-dependent noise characteristics of event cameras are used to reconstruct the intensity. However, this method was primarily developed for static scenes where there is no relative motion between the camera and the scene, which limits the advantage of the high temporal resolution of events.
@@
Recent progress in combining active illumination with event cameras [28, 29, 30, 31, 32], has resulted in accurate and high-speed geometry estimation. The setup consists of a laser scanning projector with an event camera in a stereo camera configuration. The projector illuminates the scene with a known pattern and the event camera observes the reflection of this pattern, which is then used to triangulate the depth of the scene. While it was shown [30] that the events generated by this reflection are independent of scene reflectivity, this is only true for an ideal sensor. The second order noise characteristics of an event camera, however, are illumination dependent thus creating non-idealities. This principle was used in [33] to reconstruct the absolute intensity by observing how many events were generated through the reflected light. For darker objects, the event count would be lower as the incoming light has a lower intensity, whereas brighter materials reflect more light and thus trigger more events. In this letter, we expand on previous work [34], utilizing the structured light projector not only for depth reconstruction, but also for multispectral sensing. By extracting both depth and the spectral properties of scenes in conjunction with the event camera, a unified, high-resolution, low-latency system is created, providing rich scene data for downstream tasks.
@@ III Event-based Multispectral and Depth Sensing
This section introduces the basics of an event camera (Section III-A) and multispectral imaging with an event camera (Section III-C).
III-A Event Camera
Event-cameras are unconventional, bio-inspired sensors that asynchronously measure changes (i.e., temporal contrast) in illumination at every pixel, at the time they occur [35, 36, 37, 38]. In particular, an event camera generates an event $e_{k}=(\mathbf{x}_{k},t_{k},p_{k})$ at time $t_{k}$ when the difference of logarithmic brightness at the same pixel $\mathbf{x}_{k}=(x_{k},y_{k})^{\top}$ reaches a contrast threshold $C$:
@@ III-A Event Camera
III-B Active illumination and event camera
@@
III-C Event-based Spectral Imaging
@@
In a typical event camera pixel, the photocurrent generated by the photodiode is amplified and then processed by a source follower circuit, which attempts to match its output voltage to the incoming signal. The speed at which the source follower can track changes in the signal depends on both its bias current and the amount of incoming light. Consequently, pixels exposed to brighter light respond more quickly, while those in darker regions respond more slowly [40]. Fig. 2 showcases the effect of the source follower bandwidth on the amplitude of the measured signal. This relationship between the light intensity and the source-follower voltage introduces a dependence on the scene reflectivity. Our method uses the fact that the bandwidth of the source follower in each pixel depends on the absolute intensity that its photoreceptor receives. As a result, when the illumination changes rapidly, pixels corresponding to darker or less reflective regions may not respond quickly enough to reach the event-generation threshold, even if the overall change in light intensity is significant. This means that only pixels with reflectivity above a contrast threshold will generate events, while those below this threshold will not. By carefully varying the threshold and observing which pixels generate events, we can estimate a lower bound for the reflectivity at each pixel. This method allows us to exploit the sensor’s non-idealities to recover information about the relative reflectivity of the scene.
@@ IV Characterization
This section evaluates the performance of our event-based depth and multispectral imaging system. We begin by describing the hardware setup and the evaluation baselines and performance metrics. We evaluate the performance of our system for depth estimation, comparing both qualitative and quantitative results to other depth sensors. We then evaluate the performance of our system for multispectral imaging, comparing both qualitative and quantitative results with those of other multispectral imaging systems. Finally, we evaluate the capability of our system for two tasks: material differentiation using a full-spectrum light source and material segmentation using a limited-spectrum projector and depth. To the best of our knowledge, there exists no dataset on which the proposed approach can be evaluated. Therefore, we collect our own dataset by first building our prototype system using an event camera and a projector (Fig. 3).
@@ IV-A Hardware setup
Figure 3: Our portable RGB-only setup (similar to [34]): The illumination source is a 3 wavelength laser point scanning projector (right, Sony MP-CL1A) in a stereo-camera setup with an event camera (left, Prophesee Gen3 [38]).
Our setup combines a depth camera and a projector.
We use a Prophesee Gen3 event camera [38] with a resolution of $640\times 480$ pixels. This sensor provides regular events without exposure measurements; these events are then used for depth estimation and spectral imaging. A lens with a field of view of $60^{\circ}$ is used throughout all experiments.
Figure 4: The full spectrum lab illumination setup with our event-camera seteup.
@@
The portable setup employs a Sony MP-CL1A projector with a resolution of $1920\times 720$ pixels, capable of projecting three distinct wavelengths (red, green, blue). This setup is used for data capture for color image reconstruction and the outdoor experiments. For multispectral analysis, a custom multispectral projector is built. This setup is used to compare the spectral response of our setup to a ground-truth spectrometer and a commercial multispectral camera. It consists of a full spectrum light source and six filters corresponding to wavelength bands of 10 nm each. It is combined with an optical chopper that flickers the illumination at a frequency of $100Hz$. Although highly accurate and fully customizable to any desired wavelength, this setup is too large to be used outside of the lab setting.
@@ IV-B Evaluation
The performance of the system is evaluated for depth estimation and color image reconstruction, using the portable setup (Fig. 3).
@@ V-B Real forest material differentiation demonstration
Here we demonstrate material segmentation of branches and leaves from real-world data, captured from the forested areas of the Masoala Rainforest in Zurich Zoo using the portable, $3$-wavelength illumination system (Fig. 3). The segmentation of leaves and branches can provide indispensable information for robots navigating in the forest, for instance to differentiate between pliable and traversable foliage and rigid branches to avoid. To quantitatively evaluate our approach, an automatic data labeling tool based on the Segment Anything model [45] is used. $31$ RGB-D scenes are divided into $21$ training sequences and $10$ test sequences. For each pixel, $3$ labels are predicted: (1) leaves, (2) branches or (3) background.
@@ VI Conclusion
Despite these advantages, our prototype suffers from two main limitations: First, the current portable prototype is limited to RGB multispectral sensing due to the availability of commercial projectors. Extending the system to support true multispectral sensing will require either a full-spectrum light source with appropriate wavelength filters or separate light sources, such as multispectral LEDs, for each desired wavelength. Second, the projected illumination may be insufficient under strong ambient light or if the sensor is far away from the sample, particularly for outdoor applications. Addressing this limitation will require higher output power or more efficient light sources to ensure reliable detection. Lastly, the current system is still a prototype; for proper integration into a UAV-mountable platform, the system should be redesigned as a more robust and compact sensing system.

### v2

@@ Abstract
Uncrewed aerial vehicles (UAVs) are increasingly deployed in forest environments for tasks such as environmental monitoring and search and rescue, which require safe navigation through dense foliage and precise data collection. Traditional sensing approaches, including passive multispectral and RGB imaging, suffer from latency, poor depth resolution, and strong dependence on ambient light—especially under forest canopies. In this work, we present a novel event spectroscopy system that simultaneously enables high-resolution, low-latency depth reconstruction with integrated multispectral imaging using a single sensor. Depth is reconstructed using structured light, and by modulating the wavelength of the projected structured light, our system captures spectral information in controlled bands between $650$ nm and $850$ nm. We demonstrate up to 60% improvement in RMSE over commercial depth sensors and validate the spectral accuracy against a reference spectrometer and commercial multispectral cameras, demonstrating comparable performance. A portable version limited to RGB is used to collect real-world depth and spectral data from a Masoala Rainforest. We demonstrate color image reconstruction and material differentiation between leaves and branches using this spectral and depth data. Our results show that adding depth (available at no extra effort with our setup) to material differentiation improves the accuracy by over 30% compared to color-only method. Our system, tested in both lab and real-world rainforest environments, shows strong performance in depth estimation, RGB reconstruction, and material differentiation—paving the way for lightweight, integrated, and robust UAV perception and data collection in complex natural environments.
@@ I Introduction
Figure 1: Event Spectroscopy: We propose an all-in-one solution for depth sensing, color image reconstruction and multispectral sensing. Usable for instance, for an uncrewed aerial vehicle (UAV) navigating in forest environments (C). A) Our portable setup consisting of an event camera and a projector as an illumination source B) Generated depth of scene acquired using structured light, D) spectral image reconstructed using events, E) material segmentation of leaves using spectral and depth data.
There is an increased demand for uncrewed aerial vehicles (UAVs) flying in forests [1, 2], both for environmental monitoring [3, 4], and for search and rescue applications [5]. Vision is essential for obstacle sensing and data collection to enable these applications. For instance, environmental monitoring tasks such as sensor deployment [6, 7] or sample collection [8, 4, 9] require UAVs to fly in close proximity to tree branches and foliage, where fast-moving, thin structures must be detected under dynamic and often challenging lighting conditions. Flying in close proximity to trees makes interaction with foliage probable, and while it has been shown that UAVs can push aside flexible twigs and leaves while avoiding thicker branches[10], proper sensing to differentiate between branches and foliage is needed[11]. This requires not only low latency and high resolution depth sensing of thin and fine structures, but also accurate differentiation between woody branches and foliage, for instance using multispectral sensing.
@@
In this work, we propose a single event-based structured light solution which can simultaneously deliver high-resolution and low latency depth reconstruction with integrated multispectral sensing. Based on previous work, the integrated system utilizes an event camera with structured light to reconstruct high-fidelity depth with low latency. By changing the wavelength of the light used for structured light we can not only reconstruct RGB images, but also perform multispectral sensing by projecting the desired wavelength - with lower latency than available multispectral sensors, using our novel method. To the best of our knowledge, this is the first integrated event-based system which provides both depth as well as multispectral sensing. First, we validate the quality of our depth reconstruction by demonstrating up to 60% average improvement in the RMSE over other commercial off-the-shelf depth sensors. Next, we demonstrate event-based multispectral sensing using wavelengths between 650 nm to 850 nm in a lab setting and compare our results to a commercial multispectral sensor and a ground truth spectrometer. Furthermore, we demonstrate RGB color image reconstruction using a portable version limited to RGB light projection, as well as material differentiation utilizing both the depth and spectral data on real-world data from a Masoala Rainforest.
@@ II Related Works
Our solution provides event-based multispectral and depth sensing. This requires both spectral imaging as well as event-based luminance recovery and geometry estimation. The following sections present current solutions to these problems as well as our methods.
@@ II-A Spectral Imaging
Spectral imaging captures images in multiple wavelength bands, with established, widespread applications. The resulting $(X,Y,\lambda)$ data cube contains both spatial $(X,Y)$ information and spectral information, $\lambda$, per wavelength. Data capture methods can be divided into two main groups: scanning and snapshot methods. Snapshot methods capture the full data cube during a single integration period of the sensor, whereas scanning methods achieve the same over multiple periods. Among scanning methods, tuned filters represent a popular option. These include a rotating filter wheel, an electro-mechanical Fabry-Ferot filter [14], liquid-crystals [15], or acousto-optic tunable filters [16].
@@
Recent developments in compressive sensing like CASSI [19] use a two-dimensional patterned grating to reconstruct all images of all wavelengths in a single shot. While the computational complexity required to decode the image presents a major bottleneck, improvements have made this method viable for high-resolution imaging in both spatial and spectral dimensions [20].
In this work, we employ an inverse variant of a tuned filter. Instead of filtering the light entering the imaging sensor, we illuminate the scene with light of a specific wavelength band, and measure the resulting change in reflected light off the scene with respect to ambient illumination. While this active approach requires more power and additional components compared to passive alternatives such as color filter arrays or learned methods, it enables simultaneous capture of both high fidelity spectral and depth information, and is more robust in variable and low-light conditions. Particularly for UAV applications, having an integrated system capable of providing depth, multispectral and RGB sensing as opposed to separate devices for each, compensates the increased power consumption of the active approach when considering payload restrictions.
II-B Event Luminance Recovery and Geometry Estimation
@@
As events naturally compress visual information, estimating the absolute intensity solely from events is a challenging task. Prior methods demonstrate reconstruction of a grayscale image up to an unknown intensity value through the use of data-driven priors [21]. Color intensity can also be recovered by using color event cameras [22], or through setups with multiple cameras and appropriate color filters [23]. Multispectral sensing for face recognition through the addition of a single infrared filter was shown in [24]. In [25], the illumination-dependent noise characteristics of event cameras are used to reconstruct the intensity, however, mainly for static scenes where there is no relative motion between the camera and the scene.
@@
Recent progress in combining active illumination with event cameras [26, 27, 28], has resulted in accurate and high-speed geometry estimation. The setup consists of a laser scanning projector with an event camera in a stereo camera configuration. The projector illuminates the scene with a known pattern and the event camera observes the reflection of this pattern, which is then used to triangulate the depth of the scene. While it was shown [26] that the events generated by this reflection are independent of scene reflectivity, this is only true for an ideal sensor. The second order noise characteristics of an event camera, however, are illumination dependent thus creating non-idealities. This principle was used in [29] to reconstruct the absolute intensity by observing how many events were generated through the reflected light. For darker objects, the event count would be lower as the incoming light has a lower intensity, whereas brighter materials reflect more light and thus trigger more events.
Several approaches for recovering only spectral information or only depth using event cameras have been proposed. For spectral information these include using the chromatic aberrations from a ball lens while physically adjusting the focal length of the event camera [30], to using a diffraction grating and rotating sweeping mirror to capture spectral information [31]. Looking only at depth, previous work has shown this can be reconstructed using point-spread function engineering [32] or coded apertures [33], but is missing spectral information. The utility of event cameras for downstream tasks such as semantic segmentation [34] have also been demonstrated. While non-event based, conventional systems can capture hyperspectral and depth data with a single camera using learned methods [35], the resulting depth has low resolution and fidelity.
In this letter, we expand on previous work [36], utilizing the structured light projector not only for depth reconstruction, but also for multispectral sensing. In contrast to previous approaches, by extracting both depth and the spectral properties of scenes in conjunction with the event camera, a unified, high-resolution, low-latency system is created, providing rich scene data for downstream tasks.
@@ III Event-based Multispectral and Depth Sensing
This section provides an overview of the system and introduces the basics of an event camera (Section III-B) and multispectral imaging with an event camera (Section III-D).
III-A System Overview
Our system captures both high-resolution depth and multispectral reflectivity using an event camera and a projector as a stereo pair. Depth is reconstructed using an event-based structured light method [28]. To generate depth, a point scanning laser projector shortly illuminates one pixel at a time, scanning the full projected image in the process. Due to the microsecond temporal resolution of event cameras, every pixel illumination generates an event and is matched to the origin in the projected image. From this, the depth of all illuminated pixels is then computed using standard stereo geometry. To estimate spectral reflectivity, we observe the repeated light flashes of the projector while progressively increasing the contrast threshold of the camera (reducing sensitivity). For every pixel, the minimum sensitivity at which an event still occurs indicates its relative reflectivity at that wavelength. Repeating this procedure for each spectral channel yields a multispectral image that is aligned with the depth map.
III-B Event Camera
Event-cameras are bio-inspired sensors that asynchronously measure changes (i.e., temporal contrast) in illumination at every pixel, at the time they occur [37, 38, 39, 40]. In particular, an event camera generates an event $e_{k}=(\mathbf{x}_{k},t_{k},p_{k})$ at time $t_{k}$ when the difference of logarithmic brightness at the same pixel $\mathbf{x}_{k}=(x_{k},y_{k})^{\top}$ reaches a contrast threshold $C$:
@@ III-B Event Camera
III-C Active Illumination with Event Cameras
@@
III-D Event-based Spectral Imaging
@@
In a typical event camera pixel, the photocurrent generated by the photodiode is amplified and then processed by a source follower circuit, which attempts to match its output voltage to the incoming signal. The speed at which the source follower can track changes in the signal depends on both its bias current and the amount of incoming light. Consequently, pixels exposed to brighter light respond more quickly, while those in darker regions respond more slowly [42]. Fig. 2 showcases the effect of the source follower bandwidth on the amplitude of the measured signal. This relationship between the light intensity and the source-follower voltage introduces a dependence on the scene reflectivity. Our method uses the fact that the bandwidth of the source follower in each pixel depends on the absolute intensity that its photoreceptor receives. As a result, when the illumination changes rapidly, pixels corresponding to darker or less reflective regions may not respond quickly enough to reach the event-generation threshold, even if the relative change in light intensity is significant. This means that only pixels with reflectivity above a certain threshold will generate events, while those below this threshold will not. By carefully varying the threshold and observing which pixels generate events, we can estimate a lower bound for the reflectivity at each pixel. This method allows us to exploit the sensor’s non-idealities to recover information about the relative reflectivity of the scene.
@@ IV Characterization
This section evaluates the performance of our event-based depth and multispectral imaging system. We begin by describing the hardware setup and the evaluation baselines and performance metrics. We evaluate the performance of our system for depth estimation, comparing both qualitative and quantitative results to other depth sensors. We then evaluate the performance of our system for multispectral imaging, comparing both qualitative and quantitative results with those of other multispectral imaging systems. Finally, we evaluate the capability of our system for two tasks: material differentiation using a full-spectrum light source and material segmentation using a limited-spectrum projector and depth. To the best of our knowledge, there exists no dataset on which the proposed approach can be evaluated. Therefore, we collect our own dataset by first building our prototype system using an event camera and a projector (Fig. 1A).
@@ IV-A Hardware Setup
We use a Prophesee Gen3 event camera [40] with a resolution of $640\times 480$ pixels. This sensor provides regular events without exposure measurements; these events are then used for depth estimation and spectral imaging. A lens with a field of view of $60^{\circ}$ is used throughout all experiments. The camera offers two parameters to adjust: the contrast threshold and the source follower gain. Both have the same (though inverted) downstream effect on event generation. Empirically, the working range was determined to lie between 1550 and 1800 for the contrast threshold and 415 to 600 for the source follower gain. Thus, we have a slightly higher resolution stepping of 185 for the source follow vs. 150 for the contrast threshold.
Figure 3: The full spectrum lab illumination setup with our event-camera setup.
@@
The portable setup employs a Sony MP-CL1A projector with a resolution of $1920\times 720$ pixels, capable of projecting three distinct wavelengths (red, green, blue). This setup is used for data capture for color image reconstruction and the outdoor experiments. The generation of a full depth image takes slightly less than 16 ms, and around one second to capture spectral information across the full image for one wavelength. The quality and maximum range of the depth and spectral sensing depend on the power and resolution of the projector. Most of the following scenes were captured within 1 m of the camera. For multispectral analysis, a custom multispectral projector is built (Fig. 3). This setup is used to compare the spectral response of our setup to a ground-truth spectrometer and a commercial multispectral camera. It consists of a full spectrum light source and six filters corresponding to wavelength bands of 10 nm each. It is combined with an optical chopper that flickers the illumination at a frequency of $100Hz$. Although highly accurate and fully customizable to any desired wavelength, this setup is too large to be used outside of the lab setting.
@@ IV-B Evaluation
The performance of the system is evaluated for depth estimation and color image reconstruction, using the portable setup (Fig. 1A).
@@ V-B Real Forest Material Differentiation Demonstration
Here, we demonstrate material segmentation of branches and leaves from real-world data, captured from the forested areas of the Masoala Rainforest in Zurich Zoo using the portable, $3$-wavelength illumination system (Fig. 1A). A semi-supervised segmentation pipeline is used. First, segments are generated using a min-cut algorithm, where neighboring pixels are separated if they have low similarity, based on their RGB-D vector. The segments are then classified with a shallow VGG-inspired Convolutional Neural Network with two convolutional and three fully connected layers. For classifier training, $31$ RGB-D scenes are divided into $21$ training sequences and $10$ test sequences. At the end, the pipeline predicts one of three labels for each pixel: leaves, branches or background.
@@ VI Conclusion
Our current prototype still suffers from two main limitations: the portable prototype is limited to RGB multispectral sensing due to the availability of commercial projectors. Extending the system to support true multispectral sensing will require either a full-spectrum light source with appropriate wavelength filters or separate light sources, such as multispectral LEDs, for each desired wavelength. Since every wavelength is captured separately, capture time increases in proportion to the number of wavelengths. Reducing the spatial resolution and alternating wavelengths per pixel, or projecting complimentary, non-overlapping patterns to capture multiple wavelengths simultaneously could alleviate these issues. Second, the projected illumination may be insufficient under strong ambient light or if the sensor is far away from the sample, particularly for outdoor applications. Addressing this limitation will require higher output power or more efficient light sources to ensure reliable detection. Additional inputs, for instance, polarization cues, could also improve the material classification. Lastly, the current system is still a prototype; for proper integration into a UAV-mountable platform, the system should be redesigned as a more robust and compact sensing system.